OpenStudy (anonymous):

Graph the two equations to solve the system. y = 2x -4 2x + 3y = -12

OpenStudy (zzr0ck3r):

solution is (0,-4)

OpenStudy (anonymous):

graph 3x + 2y = 6 and y = -3/2x - 1

OpenStudy (zzr0ck3r):

do you know how to graph them?

OpenStudy (anonymous):

yea i just dont know the points for this one

OpenStudy (zzr0ck3r):

the second one is y = -4 - (2/3)x

OpenStudy (anonymous):

im talkn bout to this one graph 3x + 2y = 6 and y = -3/2x - 1.. just dont know the points

OpenStudy (zzr0ck3r):

im ocnfused are we talking about another problem? 3x+2y = 6 is not in the question asked

OpenStudy (anonymous):

yea we are talkn about anotha one

OpenStudy (zzr0ck3r):

3x+2y = 6 2y = 6 - 3x y = 3 - (3/2)x so this has a t-int of 3 and a slope of -(3/2) the other one y = -(3/2)x-1 has a slope of -(2/3) and a y-int of -1, because they have the same slope and they dont have the same intercept they will never cross.

OpenStudy (zzr0ck3r):

t-int should be y-int

OpenStudy (anonymous):

so wats dah answer

OpenStudy (zzr0ck3r):

no solution

OpenStudy (anonymous):

ooo thanx

OpenStudy (zzr0ck3r):

3-(3/2)x = -1-(3/2)x 3 = -1....nope thus no solutions

OpenStudy (anonymous):

is there no solution to this y = 2 and y = -5